OpenStudy (anonymous):

Find 3 consecutive odd integers such that when the 1st is subtracted from 4 times the 3rd, the result is 7.

OpenStudy (barrycarter):

Well, if x is the first of the three integers, what are the other two (in terms of x)

OpenStudy (anonymous):

x+2 and X+3, i never now how to put this into an equation to solve, thats my only problem

OpenStudy (barrycarter):

Well, the second one is x+2, but the third one is x+4 (because odd numbers skip by 2)

OpenStudy (anonymous):

*4 sorry yeah

OpenStudy (barrycarter):

OK, so the first subtracted from 4 times the 3rd would look like what?

OpenStudy (anonymous):

x-4(x+4)

OpenStudy (barrycarter):

Almost, but it's 4(x+4)-x (since you're subtracting the first away from the 4th)

OpenStudy (anonymous):

Okay so the answer would be x=23/3

OpenStudy (barrycarter):

I'm just going to go out on a limb here and point out that 23/3 isn't an integer

OpenStudy (anonymous):

ohhhhh....

OpenStudy (anonymous):

yeah....lol

OpenStudy (anonymous):

I wrote the whole problem like this : 4(x+4)-x=7

OpenStudy (anonymous):

Distributed the 4... that gets me to, 4x+16-x+7?

OpenStudy (barrycarter):

It's sort of a trick... the answer is going to be negative

OpenStudy (anonymous):

*=7

OpenStudy (barrycarter):

Did you mean 4x+16-x=7 ?

OpenStudy (anonymous):

yes

OpenStudy (barrycarter):

OK, keep going

OpenStudy (anonymous):

4x-x=3x so 3x+16=7

OpenStudy (barrycarter):

Yup

OpenStudy (anonymous):

subtract the 16 next?

OpenStudy (barrycarter):

Good plan

OpenStudy (anonymous):

oohh -9, that gets me -3.

OpenStudy (barrycarter):

Right! You should check your answer, but the three integers are -3, what, and what?

OpenStudy (anonymous):

-1, -3, -5?

OpenStudy (barrycarter):

Ummm, -3 is the first of the 3, so not quite

OpenStudy (anonymous):

or -3,-5,and-7

OpenStudy (barrycarter):

Well, remember, the second is x+2 not x-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where those right?

OpenStudy (barrycarter):

Nope, because -5 is -3-2, not -3+2

OpenStudy (anonymous):

the answer is :5,7,9

OpenStudy (barrycarter):

@oholmes Do you have another guess/answer?

OpenStudy (anonymous):

k:integer so we have 3 odd integers:2k+1,2k+3,2k+5 2k+5-4(2k+1)=7 so k=2 that's why the 3 odd integers are:5,7,9

OpenStudy (barrycarter):

@echo_ed so 4*9 - 5 = 7 ???

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

sorry

OpenStudy (barrycarter):

OK, @oholmes did most of the work here, so the answer is: -3, -1 and +1